# Beyond Numeric Awards: In-Context Dueling Bandits with LLM Agents

### v1

@@ Abstract
In-context decision-making is an important capability of artificial general intelligence, which Large Language Models (LLMs) have effectively demonstrated in various scenarios. However, LLMs often face challenges when dealing with numerical contexts, and limited attention has been paid to evaluating their performance through preference feedback generated by the environment. This paper investigates the performance of LLMs as decision-makers in the context of Dueling Bandits (DB). We first evaluate the performance of LLMs by comparing GPT-3.5 Turbo, GPT-4, and GPT-4 Turbo against established DB algorithms. Our results reveal that LLMs, particularly GPT-4 Turbo, quickly identify the Condorcet winner, thus outperforming existing state-of-the-art algorithms in terms of weak regret. Nevertheless, LLMs struggle to converge even when explicitly prompted to do so, and are sensitive to prompt variations. To overcome these issues, we introduce an LLM-augmented algorithm, IF-Enhanced LLM, which takes advantage of both in-context decision-making capabilities of LLMs and theoretical guarantees inherited from classic DB algorithms. The design of such an algorithm sheds light on how to enhance trustworthiness for LLMs used in decision-making tasks where performance robustness matters. We show that IF-Enhanced LLM has theoretical guarantees on both weak and strong regret. Our experimental results validate that IF-Enhanced LLM is robust even with noisy and adversarial prompts.
@@ 1 Introduction
In this paper, we are particularly interested in studying the ability of Large Language Model (LLM) agent to behave optimally in online decision-making settings, i.e., have low regret. In such settings, notable failure cases have been reported [1, 2], e.g., LLM agents being vulnerable to adversarial loss functions and suffer from high regret compared to classic algorithms such as Follow-The-Regularized-Leader (FTRL) [1], and failures in exploration for classic Multi-Armed Bandit (MAB) problems [2]. These results suggest that carefully designed prompts (particularly with external summarization) and non-trivial algorithmic interventions are needed to elicit desirable behavior of LLM agents.
The failure cases encountered by LLMs may be attributed to intrinsic difficulties of handling numerical rewards, and there has been a notable lack of emphasis on evaluating the relative comparisons among the decisions they generate. To disentangle the complexities introduced by numerical rewards, this paper focuses on the problem of dueling bandits (DB) [3] as an important variant of the MAB problem. The problem of DB [3, 4] extends the classic MAB model [5] by querying for preference feedback between selected pairs of arms to identify the best one. In DB, a decision-maker chooses a pair of arms every time, but only observes the binary outcome (win or lose) of a noisy comparison between the two selected arms. Figure 1 shows a toy example illustrating the in-context interaction between an LLM agent and an environment. As a variant of the standard MAB problem, DB has attracted significant attention in recent years due to its applicability in information retrieval [6], recommendation systems [7], and online ranker evaluation [4]. This setup is particularly useful when eliciting explicit feedback is challenging or when the feedback is inherently comparative, like taste of food and product attractiveness [3]. We frame our investigation with the following question:
@@
We highlight two aspects where the capabilities and limitations of LLMs intersect with the challenges of the dueling bandit problem. First, dueling bandits pose distinctive challenges as a decision-making instance, notably due to the sparse nature of the relative rewards they provide. This sparsity complicates the in-context decision-making process, as it restricts the feedback obtained from interactions, introducing a level of difficulty not typically seen in conventional bandit problems. Even though reduction from DB to standard MAB exists [8, 9], it is unclear how LLMs would perform in DB with relative comparison feedback rather than numerical rewards. There are conceptual differences between them, similar to those for Reinforcement Learning from Human Feedback (RLHF) [10] and standard RL, where impossibility results can be found in [11]. Additionally, LLMs often struggle with numerical contexts, and there has been insufficient focus on assessing their performance using relative comparison feedback. Besides these aspects, it is worth noting that dueling bandits have broad applications, notably in areas such as healthcare decision support[12], legal judgment prediction [13], and stock movement prediction [14]. We summarize our main results below.
@@
Evaluation of LLMs’ capabilities for preference feedback. First, going beyond numeric rewards, in Section 4, we thoroughly evaluate the performance of LLM agents in terms of both strong and weak regret for making decisions in DB by comparing against various baseline DB algorithms via a case study. LLMs exhibit behaviors that significantly differ from those in classic MAB settings [2], as summarized in Table 1. Notably, GPT-4 Turbo can serve as an effective decision-maker for dueling bandits in the short term, quickly identifying and exploiting the best arm with low variance across different instances. However, the long-term performance is hindered by over-estimation bias in the exploration stage and lack of convergence criterion in the exploitation stage. This highlights the need for more robust and effective strategies to use LLMs.
Effective and Robust LLM-augmented DB algorithm. To enhance the trustworthiness of integrating LLMs in decision-making problems where performance robustness is critical, in Section 5.1, we propose a design of LLM-augmented algorithm, IF-Enhanced LLM that combines a classic DB algorithm, Interleaved Filter2 (IF2), with LLM agents. We show that the proposed algorithm has theoretical guarantees, with experiments demonstrating its efficacy and robustness with various prompting scenarios.
Figure 1: In-context interaction at time $t$ between an LLM agent and a dueling bandit environment.
@@ 2 Preliminaries
In this section, we briefly introduce the problem of dueling bandits (DB) and establish the necessary notation for this paper. Additional useful definitions can be found in Appendix A.2.1.
Dueling Bandits. In a $K$-armed dueling bandits problem, a learner interacts with the environment by selecting two arms $\mathsf{Arm}_{1}(t)$ and $\mathsf{Arm}_{2}(t)$ from a set of $K$ arms $\{b_{1},\ldots,b_{K}\}$ for a noisy comparison (a duel), at each round $t\in\{1,\ldots,T\}$ as Figure 1 illustrates. The outcome of a duel between two arms $(i,j)$ is probabilistic. More precisely, the event that an arm $b_{i}$ wins against $b_{j}$ is a Bernoulli random variable with a parameter denoted by $\Pr(b_{i}\succ b_{j})$. For notational convenience, we normalize $\Pr(b_{i}\succ b_{j})$ such that $\Pr(b_{i}\succ b_{j})=\epsilon(b_{i},b_{j})+1/2$, where $\epsilon_{ij}\coloneq\epsilon(b_{i},b_{j})\in(-1/2,1/2)$ is a measure of the distinguishability between arms $b_{i}$ and $b_{j}$, which is stationary over time and is symmetric such that $\epsilon_{ij}=-\epsilon_{ji}$ for all $i,j\in[K]\coloneq\{1,\ldots,K\}$ (see [3]). Finally, for notational convenience, we define a preference matrix $P=\left[\epsilon_{ij}\right]_{i,j\in[K]}$.
LLM Agents for Dueling Bandits. We consider an LLM agent with policy $\pi_{\mathrm{LLM}}$ interacting with a $K$-armed dueling bandit environment. At each round $t\in\{1,\ldots,T\}$, the LLM agent selects a pair of arms $(\mathsf{Arm}_{1}(t),\mathsf{Arm}_{2}(t))$ from the set $\{b_{1},\ldots,b_{K}\}$ based on a natural language instruction $\mathtt{Prompt}(C,H_{t},R)$ (see Figure 6), consisting of three parts:
  • Context $C$: a natural language description of the $K$-armed dueling bandit problem, including the number of arms $K$, the time horizon $T$, and the task objective.
  • History $H_{t}$: an externally summarized interaction history up to round $t$, which includes a sequence of pairwise dueling results and the empirical probabilities.
  • Reasoning $R$: the zero-shot chain-of-thought (CoT) reasoning [15] that encourages the LLM agent to reason about the problem in a structured manner.
@@
$\left(\mathsf{Arm}_{1}(t),\mathsf{Arm}_{2}(t)\right)=\pi_{\mathrm{LLM}}\left(\mathtt{Prompt}(C,H_{t},R)\right).$ (1)
@@
Strong and Weak Regret. We assume the standard setting that a Condorcet winner exists [7, 16, 4, 3]. The Condorcet winner $b^{*}$ is an arm that is preferred over all the other arms, i.e., $b^{*}=b_{i}$ if $\epsilon_{ij}>1/2$ for all $j\in[K]\backslash\{i\}$. Throughout this paper, we consider two important performance metrics. The first is the strong regret of a given algorithm ALG, defined as
which only compares $b^{*}$ against the better of the two selected arms $\mathsf{Arm}_{1}(t)$ and $\mathsf{Arm}_{2}(t)$. It is worth highlighting that LLM agents exhibit significantly different behaviors with respect to the two defined notions of regret, as detailed in Section 4.2. More useful definitions and settings are provided in Appendix A.2.1.
3 Related Works
Dueling Bandits. The problem of dueling bandits was initially introduced in [3]. Various methods have been proposed to tackle the task since then. These methods can be broadly classified into two categories as Explore-Then-Exploit methods and Ongoing Regret Minimization methods according to [4]. Explore-Then-Exploit methods focus on identifying the best arm with high confidence before exploiting it, such as Interleaved Filter (IF[3] and Beat the Mean (BTM[17], etc. In contrast, Ongoing Regret Minimization methods explicitly target the objective of minimizing cumulative regret, including Relative Upper Confidence Bound (RUCB[18] and Self-Sparring [7], etc. Dueling bandit problem and preference feedback in general has a wide variety of applications, including recommendation systems [3], robotics [19], and most recently, the training algorithm of large language models, such as Reinforcement Learning from Human Feedback (RLHF) [20].
LLM Agents for Multi-Armed Bandits. Several recent works have explored evaluating the capabilities of LLMs in bandit problems. For example, [21] proposed an approach to enhance contextual bandits by integrating LLMs as encoders. The LLMs’ ability to capture rich semantic and syntactic information from textual contexts is leveraged to provide the algorithm with a more informative representation of the context. The LLM-augmented algorithm transforms the raw context into a latent space vector using the LLM’s encoding capabilities. This encoded context is then used to guide the decision-making process. [2] investigates whether LLMs can engage in exploration in simple MAB environments without additional training. They compared various prompt designs and found that GPT-4 with zero-shot chain-of-thought (CoT) reasoning and an externally summarized interaction history performed the best, while other configurations failed in exploration, either by never selecting the best arm after initial rounds or by selecting all arms nearly equally often [2]. Different from the previous results, in this work we go beyond the settings of numeric rewards and investigate the capabilities of LLMs under preference feedback.
In-Context LLMs for Decision-Making. Beyond bandit problems, LLM agents have demonstrated strong capabilities in complex reasoning across a wide range of in-context decision-making tasks [22, 23, 24, 25]. Various existing works aim to understand LLM agents’ capabilities for in-context decision-making, with notable examples including planning [26, 27]. Additionally, LLM agents have been shown to enhance embodied agents in various robotic applications by providing advanced task planning abilities [28] and reward designing [29], further enabling the development of lifelong learning agents [30]. Besides these empirical successes, the authors of [1] analyzed LLMs’ interactions in online learning and game theory settings through the lens of the regret metrics. They identified simple cases where LLMs fail to be no-regret. Another line of research incorporates LLMs into classic decision-making frameworks to create LLM-augmented online decision-makers. For instance, Liu et al. [31] utilized LLMs to enhance the components of warm starting, sampling candidates, and surrogate modeling in Bayesian optimization. Our work contributes to this broad area by integrating LLM agents with the Interleaved Filter algorithm to enhance the utilization of preference feedback.
@@ 4 LLMs as Standalone In-Context Decision-Makers
To evaluate the efficacy of LLMs for solving DB problems, in this section, we use LLMs as standalone in-context decision-making agents and compare them with classic baseline algorithms. Our evaluation is two-fold: First, in Figures 2 and 8, we compare the performance of LLM agents and classic algorithms in terms of the strong and weak regret (see Eq. (2) and Eq. (3), with standard deviation). Second, we delve into the experimental results and analyze the advantages and disadvantages of LLM agents.
@@ 4.1 Implementation Details of Experiments
Prompts and Configurations of LLMs. We employ an interactive zero-shot chain-of-thought (CoT) prompt $\mathtt{Prompt}(C,H_{t},R)$, as defined in Section 2, which describes the problem setting $C$ along with an externally summarized interaction history $H_{t}$ and reasoning instructions, denoted by $R$. We adopt the prompting template that leads to the best performance among all prompt variations explored in a recent study [2] for the classic MAB problem. The LLM agents interact with dueling bandit environments in a round-based manner, with the prompt guiding their decision-making process. We conduct experiments with three LLMs: GPT-3.5 Turbo, GPT-4, and GPT-4 Turbo through the OpenAI API with temperature = 0. The detailed prompt is provided in Appendix B.1.1.
Baselines. We compare LLMs against eight well-established baseline algorithms to evaluate their efficacy. The baselines include Interleaved Filter (IF2[3], Beat the Mean (BTM[17], Sensitivity Analysis of VAriables for Generic Exploration (SAVAGE[32], Relative Upper Confidence Bound (RUCB[18], Relative Confidence Sampling (RCS[4], Relative Minimum Empirical Divergence (RMED[33], Self-Sparring [7], and Double Thompson Sampling (DTS[16]. Each of these algorithms employs distinct strategies for selecting arms and estimating preferences, with the ultimate goal of efficiently identifying the Condorcet winner. We assess the performance of LLMs and baseline algorithms using strong regret and weak regret metrics defined in Section 2.
Environments. We evaluate the regret performance of LLMs and baselines across two stochastic environments, each characterized by a distinct preference matrix $P$. The preference matrices are constructed using the Bradley-Terry-Luce (BTL) model [34, 3], with a generalized form known as the Plackett-Luce model [35]. In this model, each arm is associated with a utility parameter $\theta(i)>0$, where $i$ represents the rank of the arm (i.e., $\theta(1)$ corresponds to the best arm, $\theta(2)$ corresponds to the second best arm, and so on). For any pair of arms $b_{i}$ and $b_{j}$, the probability of $b_{i}$ being preferred over $b_{j}$ is determined by $P{(i\succ j)}=\theta(i)/(\theta(i)+\theta(j))$. Setting the number of arms $K=5$, we randomize the order of the arms to prevent selection bias, resulting in the following arm ordering: $b_{5}\succ b_{3}\succ b_{2}\succ b_{1}\succ b_{4}$. We use two instances: Easy and Hard, with their respective $\theta$ parameters given by:
  • Easy instance: $\theta(1)=1,\ \theta(i)=0.5-(i-1)/2K$,  $\forall i\in[2,K]$.
  • Hard instance: $\theta(i)=1-(i-1)/K$,  $\forall i\in[K]$.
Note that the datasets generated in this way satisfy the Strong Stochastic Transitivity (SST) and Stochastic Triangle Inequality (STI) properties [3] (see Appendix A.2.1 for more details). The settings of the BTL model used in our experiments also imply the existence of a Condorcet winner.
@@
Figure 2: Comparisons between LLM agents and various classic DB algorithms. Left and Right: strong and weak regret for the Easy instance. Results for the Hard instance are presented in Figure 8.

### v3

@@ Abstract
In-context reinforcement learning (ICRL) is a frontier paradigm for solving reinforcement learning problems in the foundation model era. While ICRL capabilities have been demonstrated in transformers through task-specific training, the potential of Large Language Models (LLMs) out-of-the-box remains largely unexplored. Recent findings highlight that LLMs often face challenges when dealing with numerical contexts, and limited attention has been paid to evaluating their performance through preference feedback generated by the environment. This paper is the first to investigate LLMs as in-context decision-makers under the problem of Dueling Bandits (DB), a stateless preference-based reinforcement learning setting that extends the classic Multi-Armed Bandit (MAB) model by querying for preference feedback. We compare GPT-3.5 Turbo, GPT-4, GPT-4 Turbo, Llama 3.1, and o1-preview against nine well-established DB algorithms. Our results reveal that our top-performing LLM, GPT-4 Turbo, has the zero-shot relative decision-making ability to achieve surprisingly low short-term weak regret across all the DB environment instances by quickly including the best arm in duels. However, an optimality gap exists between LLMs and classic DB algorithms in terms of strong regret. LLMs struggle to converge and consistently exploit even when explicitly prompted to do so, and are sensitive to prompt variations. To bridge this gap, we propose an agentic flow framework: LLM with Enhanced Algorithmic Dueling (LEAD), which integrates off-the-shelf DB algorithms with LLM agents through fine-grained adaptive interplay. We show that LEAD has theoretical guarantees inherited from classic DB algorithms on both weak and strong regret. We validate its efficacy and robustness even with noisy and adversarial prompts. The design of such an agentic framework sheds light on how to enhance the trustworthiness of general-purpose LLMs used for in-context decision-making tasks.
@@ 1 Introduction
Large sequence models pretrained with task specific interactive datasets have led to the emergence of in-context reinforcement learning (ICRL) (Laskin et al., 2022; Lee et al., 2024), where models can infer tasks from interaction histories as context and make effective decisions in unseen environments without parameter updates. Through trial and error, these models can self-improve their policies purely in-context. While ICRL capabilities have been demonstrated in transformers with task-specific training from scratch, the potential of general-purpose Large Language Models (LLMs) to perform ICRL remains largely unexplored. Recent investigations into LLMs’ out-of-the-box ICRL capabilities in environments with numeric rewards have reported notable failure cases, e.g., LLM agents being vulnerable to adversarial loss functions and suffering from high regret compared to classic algorithms such as Follow-The-Regularized-Leader (FTRL) (Park et al., 2024), and exhibiting failures in exploration within Multi-Armed Bandit (MAB) problems via standard training (Krishnamurthy et al., 2024). Even with inference-time algorithmic guidance, an optimality gap persists between LLMs and classic (contextual) MAB algorithms (Nie et al., 2024). These results suggest that carefully designed prompts and non-trivial algorithmic interventions are needed to elicit desirable in-context reinforcement learning behavior of LLM agents.
The failure cases encountered by LLMs may be attributed to intrinsic difficulties in processing numeric rewards, especially in tasks where patterns are difficult to express in natural language. Recent findings have pointed out that LLMs often struggle with simple numerical comparisons (e.g., incorrectly judging 13.11 to be larger than 13.8), and there has been a notable lack of emphasis on evaluating the relative comparisons among the decisions they generate. Figure 1 shows a toy example illustrating the in-context interaction between an LLM agent and different environment settings. To disentangle the complexities introduced by numerical rewards, this paper focuses on the problem of Dueling Bandits (DB) (Yue et al., 2012; Zoghi et al., 2014a), a stateless preference-based reinforcement learning setting (Wirth et al., 2017; Pacchiano et al., 2021) that extends the classic MAB model by querying for preference feedback between selected pairs of arms to identify the best one. In DB, the agent learns through binary outcome (win or lose) of a noisy comparison between the two selected arms. This setup is particularly useful when eliciting explicit feedback is challenging or when the feedback is inherently comparative, like taste of food and product attractiveness (Yue et al., 2012). DB has attracted significant attention due to its applicability in information retrieval (Yue and Joachims, 2009), recommendation systems (Sui et al., 2017), and online ranker evaluation (Zoghi et al., 2014a). We frame our investigation with the following question:
@@
The DB problem poses distinctive challenges as a relative decision-making instance, particularly due to the sparse nature of the relative rewards. This sparsity complicates the in-context decision-making process, as it restricts the feedback obtained from interactions, introducing a level of difficulty not typically seen in conventional bandit problems. Even though reduction from DB to standard MAB exists (Ailon et al., 2014; Saha and Gaillard, 2022), it remains unclear how LLMs would perform in DB with preference feedback rather than numeric rewards. There are conceptual differences between them, similar to those between Reinforcement Learning from Human Feedback (RLHF) (Stiennon et al., 2020) and standard RL, where impossibility results can be found in (Wang et al., 2024a).
While task-specific training of large sequence models can yield promising ICRL results, it is often impractical due to the substantial computational resources required. Similar to the settings in (Krishnamurthy et al., 2024; Nie et al., 2024; Mirchandani et al., 2023; Chen et al., 2024), we evaluate the emergent zero-shot abilities (Wei et al., 2022) of ICRL in general-purpose LLMs under the dueling bandit problem, without re-training or fine-tuning. We summarize our main results below.
Figure 1: In-context reinforcement learning of an LLM agent with numeric rewards (in a multi-armed bandit environment) and preference feedback (in a dueling bandit environment).
@@
Evaluation of LLMs’ Emergent Zero-Shot Abilities of In-Context Dueling Bandits. We go beyond numeric rewards to evaluate the performance of LLM agents in terms of both strong and weak regret for making decisions in DB by comparing against various baseline DB algorithms via a case study. We found that the top-performing general-purpose LLMs has the zero-shot ability for relative decision-making sufficient to achieve low weak regret in DB, which significantly differs from that in classic MAB settings (Krishnamurthy et al., 2024). Notably, GPT-4 Turbo can serve as an effective decision-maker for dueling bandits in terms of weak regret, quickly selecting the best arm in duels with low variance across a range of instances. However, consistent with (Nie et al., 2024), we found that an optimality gap exists between LLMs and classic DB algorithms in terms of strong regret. LLMs’ performance is hindered by over-estimation bias in the exploration stage and lack of convergence criterion in the exploitation stage. This highlights the need for more effective and robust strategies to bridge this gap for in-context dueling bandits.
Effective and Robust Agentic Flow Framework for In-Context Dueling Bandits. To address the identified optimality gap and enhance the trustworthiness of in-context LLM agents in DB tasks, in Section 4.1, we propose an agentic flow framework, LLM with Enhanced Algorithmic Dueling (LEAD) that integrates off-the-shelf Explore-then-Exploit DB algorithms with LLM agents. This framework enables the fine-grained adaptive interplay between rule-based expert systems and in-context LLM agents, enhancing their ability to handle DB problems via algorithmic interventions as suggested in (Krishnamurthy et al., 2024; Nie et al., 2024). As an illustrative example, we demonstrate how Interleaved Filter2 (IF2) algorithm can be incorporated with LLM agents in this framework. We show that LEAD has theoretical guarantees, with experiments demonstrating its efficacy and robustness across various prompting scenarios.
@@ 2 Preliminaries
In this section, we briefly introduce the problem of dueling bandits (DB) and establish the necessary notation for this paper. Additional useful definitions can be found in Appendix B.3.1.
Dueling Bandits. In a fundamental $K$-armed dueling bandit problem setting (Yue et al., 2012), a learner interacts with the environment by selecting two arms $\mathsf{Arm}_{1}(t)$ and $\mathsf{Arm}_{2}(t)$ from a set of $K$ arms $\{b_{1},\ldots,b_{K}\}$ for a noisy comparison (a duel), at each round $t\in\{1,\ldots,T\}$ as Figure 1 illustrates. The outcome of a duel between two arms $(i,j)$ is probabilistic. More precisely, the event that an arm $b_{i}$ wins against $b_{j}$ is a Bernoulli random variable with a parameter denoted by $\Pr(b_{i}\succ b_{j})$. For notational convenience, we normalize $\Pr(b_{i}\succ b_{j})$ such that $\Pr(b_{i}\succ b_{j})=\epsilon(b_{i},b_{j})+1/2$, where $\epsilon_{ij}\coloneq\epsilon(b_{i},b_{j})\in(-1/2,1/2)$ is a measure of the distinguishability between arms $b_{i}$ and $b_{j}$, which is stationary over time and is symmetric such that $\epsilon_{ij}=-\epsilon_{ji}$ for all $i,j\in[K]\coloneq\{1,\ldots,K\}$. Finally, for notational convenience, we define a preference matrix $P=\left[\epsilon_{ij}\right]_{i,j\in[K]}$.
In-Context LLM Agents for Dueling Bandits. We consider an LLM agent with policy $\pi_{\mathrm{LLM}}$ interacting with a $K$-armed dueling bandit environment in-context. At each round $t\in\{1,\ldots,T\}$, the LLM agent selects a pair of arms $(\mathsf{Arm}_{1}(t),\mathsf{Arm}_{2}(t))$ from the set $\{b_{1},\ldots,b_{K}\}$ based on a natural language instruction $\mathtt{Prompt}(C,H_{t},R)$ (see Figure 7), consisting of three parts:
  • Problem Description $P$: a natural language description of the DB problem, including the number of arms $K$, the time horizon $T$, and the task objective.
  • History $H_{t}$: an externally summarized interaction history (Krishnamurthy et al., 2024) up to round $t$, which includes a sequence of pairwise dueling results and the empirical probabilities.
  • Reasoning $R$: the zero-shot chain-of-thought (CoT) reasoning (Kojima et al., 2022) that encourages the LLM agent to reason about the problem in a structured manner.
@@
$\left(\mathsf{Arm}_{1}(t),\mathsf{Arm}_{2}(t)\right)=\pi_{\mathrm{LLM}}\left(\mathtt{Prompt}(P,H_{t},R)\right).$ (1)
@@
Strong and Weak Regret. Throughout this paper, we assume the standard setting that a Condorcet winner (CW) exists (Sui et al., 2017; Wu and Liu, 2016; Zoghi et al., 2014a; Yue et al., 2012). The CW denoted as $b^{*}$ is an arm that is preferred over all the other arms, i.e., $b^{*}=b_{i}$ if $\epsilon_{ij}>1/2$ for all $j\in[K]\backslash\{i\}$. We consider two performance metrics: (i) strong regret (SR), which evaluates the total preference gap between $b^{*}$ and both selected arms; (ii) weak regret (WR), which compares $b^{*}$ only with the better of the two arms. Detailed definitions and settings are provided in Appendix B.3.1.
Related Works. Our work contributes to the growing community of intersection between LLMs and decision-making. We summarize the detailed related works about dueling bandits, LLM agents for bandits, and LLMs for in-context decision-making in the Appendix A.
@@ 3 LLMs as Standalone In-Context Decision-Makers
To evaluate the LLMs’ efficacy for solving DB problems in-context, in this section, we use LLMs as standalone decision-making agents and compare them with classic DB algorithms. Our evaluation is two-fold: First, in Figures 2 and 9, we compare the performance of LLMs and classic algorithms in terms of the strong and weak regret (see Eq.(4) and Eq.(5), with standard deviation). Second, we delve into the experimental results and analyze the success and failure modes of LLM agents.
@@ 3.1 Implementation Details of Experiments
Prompts and Configurations of LLMs. We employ an interactive zero-shot chain-of-thought (CoT) prompt $\mathtt{Prompt}(P,H_{t},R)$, as defined in Section 2, which describes the problem setting $P$, externally summarized interaction history $H_{t}$ and reasoning instructions $R$. We adopt the prompting template and LLM configurations that lead to the best performance among all prompt variations explored in recent studies (Krishnamurthy et al., 2024; Nie et al., 2024) for MAB problem. The LLM agents interact with dueling bandit environments in a round-based manner, with the prompt guiding their decision-making process. We conduct experiments with five LLMs: GPT-3.5 Turbo, GPT-4, GPT-4 Turbo, Llama 3.1, and o1-preview. Note that we skip the GPT-4o version which is primarily developed for multimodal tasks and has the same intelligence as GPT-4 Turbo. The detailed prompt is provided in Appendix C.1.2.
Baselines. We compare LLMs against nine well-established baseline algorithms to evaluate their efficacy. The baselines include Interleaved Filter (IF2(Yue et al., 2012), Beat the Mean (BTM(Yue and Joachims, 2011), Sensitivity Analysis of VAriables for Generic Exploration (SAVAGE(Urvoy et al., 2013), Relative Upper Confidence Bound (RUCB(Zoghi et al., 2014b), Relative Confidence Sampling (RCS(Zoghi et al., 2014a), Relative Minimum Empirical Divergence (RMED(Komiyama et al., 2015), Versatile Dueling Bandits (VDB(Saha and Gaillard, 2022), Self-Sparring (Sui et al., 2017), and Double Thompson Sampling (DTS(Wu and Liu, 2016). Each of these algorithms employs distinct strategies for selecting arms and estimating preferences, with the ultimate goal of efficiently identifying the Condorcet winner. We assess the performance of LLMs and baseline algorithms using strong regret and weak regret metrics defined in Section 2.
[…]